# On integer distance sets

### v2

@@ Abstract.
We develop a new approach to address some classical questions concerning the size and structure of integer distance sets. Our main result is that any integer distance set in the Euclidean plane has all but a very small number of points lying on a single line or circle. From this, we deduce a near-optimal lower bound on the diameter of any non-collinear integer distance set of size $n$ and a strong upper bound on the size of any integer distance set in $[-N,N]^{2}$ with no three points on a line and no four points on a circle.
@@ 1. Introduction
In this paper, we develop a new approach to the study of integer distance sets that enables us to prove a structure theorem partially explaining this phenomenon, showing that any integer distance set in $[-N,N]^{2}$ has all but at most polylogarithmically many points lying on a single line or circle.
@@
Let $S\subset[-N,N]^{2}$ be an integer distance set. Then, there exists a line or circle $C\subset\mathbf{R}^{2}$ such that
$\left|S\setminus C\right|=O\left((\log{N})^{O(1)}\right).$
@@
Theorem 1.1 is the first-ever general structure theorem for integer distance sets, and we use it to address two classical questions about integer distance sets. The first is the question of how large an integer distance set can be if it has no three points on a line and no four points on a circle (see, for example, [10, Chapter 13]). This question is often attributed to Erdős, though he says in [13] that it is an “old problem whose origin I cannot trace”. Erdős still posed it on numerous occasions (for instance, in [12]), and, in particular, asked whether such a set consisting of seven points exists in the plane. In 2008, Kreisel and Kurz [23] indeed found seven points in the plane with no three of them on a line, no four of them on a circle, and with pairwise integer distances. No larger such set has been found since.
@@
The previous best known upper bound for the size of integer distance sets in $[-N,N]^{2}$ with no three points on a line and no four points on a circle was $O(N)$, which follows from work of Solymosi [30]. It could even be the case that there is a uniform upper bound on the size of any integer distance set with no three points on a line and no four points on a circle. Indeed, Ascher, Braune, and Turchet [2] proved this conditionally on Lang’s conjecture.
The second question we address is that of estimating the minimum diameter of a non-collinear integer distance set of size $n$ (see, for example, Problem 8 of [5, Section 5.11]). It is known, thanks to work of Solymosi [30] and Harborth, Kemnitz, and Möller [16], that this minimum diameter has order of magnitude at least $n$ and at most $n^{O(\log\log{n})}$. Theorem 1.1 reduces the problem of proving a lower bound for the diameter to the problem of proving an upper bound for the maximum number of collinear and concyclic points a non-collinear integer distance set in $[-N,N]^{2}$ can contain. The paper [24] of Kurz and Wassermann implicitly contains such a bound for the number of collinear points.
@@
Bat-Ochir [3] proved an upper bound of the same strength for circles contained in $[-N,N]^{2}$ with special radii, which we extend to all circles.
@@
Combining Theorem 1.1 with Propositions 1.4 and 1.5, we improve Solymosi’s lower bound [30] to obtain almost tight bounds on the smallest possible diameter of a non-collinear integer distance set of size $n$.
@@
It is possible to make all implied constants in the proofs of Propositions 1.4 and 1.5 completely explicit, and thus to compute an explicit $c>0$ such that the minimum diameter of a non-collinear integer distance set of size $n$ has order of magnitude at least $n^{c\log\log{n}}$. On the other hand, one can also compute, using the constructions in [16], an explicit constant $c^{\prime}>c$ such that the minimum diameter of a non-collinear integer distance set of size $n$ has order of magnitude at most $n^{c^{\prime}\log\log{n}}$. We can improve these constructions a bit to obtain a $c^{\prime}$ slightly smaller than the best known one currently in the literature, but there is still a gap between this and the largest possible value of $c$ produced by our arguments. It is, therefore, still an open problem to determine the exact asymptotics of the minimum diameter of a non-collinear integer distance set of size $n$. Kurz and Wassermann [24] conjectured that, for $n$ sufficiently large, an integer distance set of size $n$ with minimum diameter has all but one point collinear. This problem is also still open. Finally, the famous question of Erdős and Ulam of whether there exists a dense rational distance set in the plane is still unresolved, though the answer is known to be negative conditionally on the $abc$ conjecture [25] or on Lang’s conjecture [29, 31].
@@
All previous progress towards Corollaries 1.3 and 1.6 relied heavily on this argument of Erdős, which is one of many applications of “low degree algebraic geometry” to extremal combinatorics and discrete geometry. It appears that the utility of low degree algebraic geometry in understanding the size and structure of integer distance sets is limited, however, as can be seen from the lack of progress beyond Solymosi’s work [30]. Our new innovation is to use “intermediate degree algebraic geometry” to study integer distance sets; we expect this general method to have further applications in combinatorial and discrete geometry.
@@
    We show that any integer distance set $S$ in $[-N,N]^{2}$ can be transformed into a set $\widetilde{S}=LS$ of rational points of height $O(N^{2})$ via an appropriate invertible linear transformation $L$ of $\mathbf{C}^{2}$ (see Lemma 2.1 below).
@@
    According to the above analysis, we would ideally like to show that if $S$ is not almost completely contained in a single line or circle, then $\widetilde{S}$ lies in the union of $O(\log N)^{O(1)}$ irreducible algebraic curves, defined over $\mathbf{Q}$, each of degree roughly $(\log N)^{C}$ (for some constant $C$).
@@
    We prove that the set $X_{k}$ of all such $(\widetilde{x},\widetilde{y},d_{1},\ldots,d_{k})$ (as $(\widetilde{x},\widetilde{y})$ ranges over $\mathbf{C}^{2}$) is an irreducible algebraic surface, defined over $\mathbf{Q}$, of degree roughly $(\log N)^{C}$. Since $S$ is an integer distance set, the set $\widetilde{S}^{\prime}$ of all lifted points $(\widetilde{x},\widetilde{y},d_{1},\ldots,d_{k})$ of $\widetilde{S}$ on $X_{k}$ consists of rational points of height $O(N^{2})$. Therefore, we are close to the aim in (2): we have transformed our points into rational points of the correct height inside an irreducible algebraic variety of the correct degree. However, this variety is a surface, rather than a curve.
@@
    To tackle this issue, we begin by covering $\widetilde{S}^{\prime}$ by $O((\log N)^{O(1)})$ irreducible algebraic curves $\widetilde{\gamma}$, defined over $\mathbf{Q}$, of degree $O((\log N)^{C})$, contained in the surface $X_{k}$. The existence of such curves is ensured by the results of [6] (see Theorem 2.2). Projecting the first two coordinates to $\mathbf{C}^{2}$, we are now closer to the aim in (2), but not quite there: we have covered $\widetilde{S}$ by an acceptable number of irreducible algebraic curves $\widetilde{\gamma}$ defined over $\mathbf{Q}$. However, these curves may have very low degree relative to the desired $(\log N)^{C}$.
@@
    We resolve this issue by noticing that we have a lot of freedom to choose the points $P_{1},\ldots,P_{k}$ we fix at the start of the argument. Each choice leads to a different $X_{k}$, with properties as in (4). We prove that, if $L^{-1}\widetilde{\gamma}$ is not a line or a circle, or if it is but there are enough points in $S$ outside it, then there is a choice of initial points $P_{1},\ldots,P_{k}$ such that the lift of $\widetilde{\gamma}$ on the corresponding $X_{k}$ is an irreducible curve of the correct degree $(\log N)^{C}$. This completes the proof.
@@ 2.1. Encoding the points of $S$ as rational points on a surface
We will use basic definitions and results from algebraic geometry throughout this paper; a good reference for this is [18, Chapter I]. Observe that it suffices to prove Theorem 1.1 in the case that $S$ contains the origin and at least one other point. Indeed, either $S$ contains at least two points $P$ and $P^{\prime}$ in $[-N,N]^{2}$, or the result is trivial; moreover, if there exists a line or circle $C\subset\mathbf{R}^{2}$ such that all but $O((\log{2N})^{O(1)})$ points of $(S-P)\cap[-2N,2N]^{2}$ are contained in $C$, then certainly all but $O((\log{N})^{O(1)})$ points of $S$ are contained in a single line or circle. Thus, it follows from Lemma 2.1 that we may further assume, without loss of generality, that there exists a squarefree $m\in\mathbf{N}$ such that
@@
Let $k=k(N)$ be an integer parameter increasing slowly with $N$ (which we will, eventually, take to be $\asymp\log\log{N}$). If $|S|\leq k+1$, then we are done. Otherwise, pick $k$ distinct points $\tilde{P}_{j}=(a_{j},b_{j}\sqrt{m})\in S$ (so that $a_{i},b_{i}\in\frac{1}{2M}\mathbf{Z}\subseteq\mathbf{Q}$) for $1\leq j\leq k$, none of which is the origin, and set $P_{j}=(a_{j},b_{j})$ for each $1\leq j\leq k$. For any point $P=(a,b)\in\mathbf{Q}^{2}$, we define the polynomial $Q_{m,P}\in\mathbf{Q}[x,y,d]$ by
@@ 2.3. Counting rational points of small height on curves
$\mathcal{C}_{k}:=\{(x,y,d_{1},\dots,d_{k}):Q_{m,P_{j}^{\prime}}(x,y,d_{j})=0\text{ for all }j=1,\dots,k\}\cap(C^{\prime}\times\mathbf{C}^{k})\subset\mathbf{C}^{k+2}.$
Then, the points in $S\cap C$ correspond to a certain subset of rational points in the projective closure $\overline{\mathcal{C}_{k}}\subset\mathbf{P}^{k+2}$ of height at most $8N^{2}$. In Section 4, we will prove that $\overline{\mathcal{C}_{k}}$ is an irreducible curve of degree between $2^{k}$ and $2^{k}\deg{C}$, provided that $\tilde{P}_{1}^{\prime},\dots,\tilde{P}_{k}^{\prime}\in S$ are chosen appropriately. We emphasise that these points will be chosen on $C$ when $C$ is not a line or a circle, and outside of $C$ otherwise.
We can then apply a refinement of the Bombieri–Pila [4] bound due to Castryk, Cluckers, Dittmann, and Nguyen [6], who proved an upper bound on $|V(\mathbf{Q},N)|$, when $V\subset\mathbf{P}^{n}$ is an irreducible algebraic curve defined over $\mathbf{Q}$, with optimal dependence on $N$ and with an explicit polynomial dependence on the degree of $V$. By tracing through the proof of [6, Theorem 2] (and the results used to deduce it from the $n=2$ case [6, Proposition 4.3.2 and Lemma 5.1]), one can see that the dependence of their implied constant on $n$ is $\ll n^{200n^{2}/d}d^{4n^{2}/d}$. We record this in the following theorem.
@@
Thus, as long as $C$ either is not a line or circle or is a line or circle for which $|S\setminus C|\gg(\log\log{N})^{2}$, we can use Theorem 2.3 to bound the number of points in $S\cap C$ by $\ll e^{O(k)}N^{2/2^{k}}$, which, for $k\asymp\log\log N$, is $\ll(\log N)^{O(1)}$. Putting everything together yields Theorem 1.1. We will carry out the argument just outlined in Sections 34, and 6, and prove Propositions 1.4 and 1.5 in Section 5.
@@
Some polynomial dependence on $d$ is necessary in the bounds in Theorems  2.2 and 2.3 (see [6, Proposition 5]). As a consequence, no bound better than $|S\setminus C|\ll(\log{N})^{O(1)}$ in Theorem 1.1 seems attainable via our methods. The determinant method produces bounds that apply to any irreducible surface (respectively, curve) in $\mathbf{P}^{n}$ of degree $d$, but significantly better bounds could hold for individual surfaces (respectively, curves) in this collection. Thus, a further analysis of the specific varieties arising in our argument would be needed to improve Theorem 1.1.

### v3

@@ Abstract.
We develop a new approach to address some classical questions concerning the size and structure of integer distance sets. Our main result is that any integer distance set in the Euclidean plane is either very sparse or has all but an exceedingly small proportion of its points lying on a single line or circle. From this, we deduce a near-optimal lower bound on the diameter of any non-collinear integer distance set of size $n$ and a strong upper bound on the size of any integer distance set in $N,N]^{2}$ with no three points on a line and no four points on a circle.
@@ 1. Introduction
In this paper, we develop a new approach to the study of integer distance sets that enables us to prove a structure theorem partially explaining this phenomenon, showing that any integer distance set in $N,N]^{2}$ must either be polylogarithmically small or have the vast majority of its points lying on a single line or circle.
@@
Let $S\subset[-N,N]^{2}$ be an integer distance set. Then, either $\left|S\right|=O\left((\log{N})^{O(1)}\right)$, or else there exists a line or circle $C\subset\mathbf{R}^{2}$ such that
$\left|S\setminus C\right|=O\left((\log\log{N})^{2}\right).$
@@
Theorem 1.1 is the first-ever general structure theorem for integer distance sets, and we use it to address two classical questions about integer distance sets.
1.1. Erdős’s integer distance set problem
The first question we address is the question of how large an integer distance set can be if it has no three points on a line and no four points on a circle (see, for example, [11, Chapter 13]). This question is often attributed to Erdős, though he says in [14] that it is an “old problem whose origin I cannot trace”. Erdős still posed it on numerous occasions (for instance, in [13]), and, in particular, asked whether such a set consisting of seven points exists in the plane. In 2008, Kreisel and Kurz [30] indeed found seven points in the plane with no three of them on a line, no four of them on a circle, and with pairwise integer distances. No larger such set has been found since.
@@ 1.1. Erdős’s integer distance set problem
The previous best known upper bound for the size of integer distance sets in $N,N]^{2}$ with no three points on a line and no four points on a circle was $O(N)$, which follows from work of Solymosi [38]. It could even be the case that there is a uniform upper bound on the size of any integer distance set with no three points on a line and no four points on a circle. Indeed, Ascher, Braune, and Turchet [3] proved this conditionally on Lang’s conjecture.
Remark 1.4.
Interestingly, Erdős’s integer distance set problem is related to Fuglede’s problem on the maximal size of frequency sets $\Lambda\subset\mathbf{R}^{2}$ whose corresponding system
$E(\Lambda)=\left\{e^{2\pi i\lambda\cdot x}\colon\lambda\in\Lambda\right\}$
of exponential functions is orthogonal in $L^{2}(D)$, where $D=\{x\in\mathbf{R}^{2}\colon\|x\|<1\}$ is the unit disk (see [27], where this relation is explored). Fuglede proved that any such $\Lambda$ must be finite [15, 16] (see also [24, 25, 27] for related results).
Note that the system $E(\Lambda)$ is orthogonal in $L^{2}(D)$ if and only if
$\int_{D}e^{2\pi i(\lambda-\lambda^{\prime})\cdot x}dx=\widehat{\mathbf{1}}_{D}(\lambda^{\prime}-\lambda)=0$
for any pair of distinct points $\lambda,\lambda^{\prime}$ in $\Lambda$, and recall that $\xi\in\mathbf{R}^{2}$ is a zero of $\widehat{\mathbf{1}}_{D}$ if and only if $2\pi\|\xi\|$ is a zero of the Bessel function $J_{1}$. Therefore, the system $E(\Lambda)$ is orthogonal in $L^{2}(D)$ if and only if the distance set of $\Lambda$,
[…]
is contained in the zero-set of the function $J_{1}(2\pi\>\cdot)$. Using the asymptotic expansion [1] of the function $J_{1}(2\pi\>\cdot)$, we have that its positive zeros, and hence the distances (1.1), are of the form
$\frac{n}{2}+\frac{1}{8}+O\left(\frac{1}{n}\right),\quad n\in\mathbf{N}.$
Since the set $\left\{\frac{n}{2}+\frac{1}{8}\colon n\in\mathbf{Z}\right\}$ is not closed under addition, this then imposes (by, e.g., [26, Lemma 1] and Ptolemy’s theorem) geometric constraints on the frequency set $\Lambda$ that are very similar to the constraints in Erdős’s integer distance set problem, i.e., (approximate) lines and circles are excluded. Using these constraints, bounds on the size of $\Lambda\cap[-N,N]^{2}$ were established in [26, 41], with the current record of $O_{\epsilon}(N^{3/5+\epsilon})$ due to Zakharov in [41].
Fuglede conjectured [15] that any frequency set $\Lambda\subset\mathbf{R}^{2}$ that gives rise to an orthogonal system $E(\Lambda)$ in $L^{2}(D)$ cannot contain more than three points. Indeed, if $|\Lambda|\geq 4$ for such a set $\Lambda$, then there would be algebraic relations among the distances (1.1), since the $5\times 5$ Cayley–Menger determinant of any planar four-point simplex vanishes. This, in turn, would imply non-trivial algebraic relations between zeros of $J_{1}$, which contradicts the expectation that these zeros are algebraically independent. Nevertheless, since the non-existence of algebraic relations among zeros of $J_{1}$ is yet an open problem in transcendental number theory, this conjecture of Fuglede remains open.
It is plausible, however, that one could significantly improve upon Zakharov’s bound by adapting the machinery we introduce in this work, namely, encoding the points of $\Lambda$ as lattice points on an analytic manifold [^1] For instance, let $\varphi\colon\mathbb{R}\to\mathbb{R}$ be an analytic function whose integer points are the zeros of the Bessel function $J_{1}(2\pi\>\cdot)$, and let $\mathcal{CM}\colon\mathbb{R}^{6}\to\mathbb{R}$ be the $5\times 5$ Cayley–Menger determinant. Fix three distinct points $P_{1},P_{2},P_{3}\in\Lambda$. Then, for any point $P\in\Lambda\setminus\{P_{1},P_{2},P_{3}\}$, the simplex $(P,P_{1},P_{2},P_{3})$ is assigned to an integer point $(\varphi^{-1})^{\otimes 6}\left(\|P_{1}-P_{2}\|,\|P_{1}-P_{3}\|,\|P_{1}-P_{4}\|,\|P_{2}-P_{3}\|,\|P_{2}-P_{4}\|,\|P_{3}-P_{4}\|\right)$ on the manifold $\mathcal{CM}(\varphi^{\otimes 6})=0$. and then using the Pila–Wilkie theorem [34].
1.2. Sharp diameter bound
The second question we address is that of estimating the minimum diameter of a non-collinear integer distance set of size $n$ (see, for example, Problem 8 of [6, Section 5.11]). It is known, thanks to work of Solymosi [38] and Harborth, Kemnitz, and Möller [19], that this minimum diameter has order of magnitude at least $n$ and at most $n^{O(\log\log{n})}$. Theorem 1.1 reduces the problem of proving a lower bound for the diameter to the problem of proving an upper bound for the maximum number of collinear and concyclic points a non-collinear integer distance set in $N,N]^{2}$ can contain. The paper [31] of Kurz and Wassermann implicitly contains such a bound for the number of collinear points.
@@ 1.2. Sharp diameter bound
Bat-Ochir [4] proved an upper bound of the same strength for circles contained in $N,N]^{2}$ with special radii, which we extend to all circles.
@@
Combining Theorem 1.1 with Propositions 1.5 and 1.6, we improve Solymosi’s lower bound [38] to obtain almost tight bounds on the smallest possible diameter of a non-collinear integer distance set of size $n$.
@@
It is possible to make all implied constants in the proofs of Propositions 1.5 and 1.6 completely explicit, and thus to compute an explicit $c>0$ such that the minimum diameter of a non-collinear integer distance set of size $n$ has order of magnitude at least $n^{c\log\log{n}}$. On the other hand, one can also compute, using the constructions in [19], an explicit constant $c^{\prime}>c$ such that the minimum diameter of a non-collinear integer distance set of size $n$ has order of magnitude at most $n^{c^{\prime}\log\log{n}}$. We can improve these constructions a bit to obtain a $c^{\prime}$ slightly smaller than the best known one currently in the literature, but there is still a gap between this and the largest possible value of $c$ produced by our arguments. It is, therefore, still an open problem to determine the exact asymptotics of the minimum diameter of a non-collinear integer distance set of size $n$. Kurz and Wassermann [31] conjectured that, for $n$ sufficiently large, an integer distance set of size $n$ with minimum diameter has all but one point collinear. This problem is also still open. Finally, the famous question of Erdős and Ulam of whether there exists a dense rational distance set in the plane is still unresolved, though the answer is known to be negative conditionally on the abcabc conjecture [32] or on Lang’s conjecture [37, 39].
@@
All previous progress towards Corollaries 1.3 and 1.7 relied heavily on this argument of Erdős, which is one of many applications of “low degree algebraic geometry” to extremal combinatorics and discrete geometry. It appears that the utility of low degree algebraic geometry in understanding the size and structure of integer distance sets is limited, however, as can be seen from the lack of progress beyond Solymosi’s work [38]. Our new innovation is to use “intermediate degree algebraic geometry” to study integer distance sets; we expect this general method to have further applications in combinatorial and discrete geometry.
@@
    We show that any integer distance set $S$ in $N,N]^{2}$ can be transformed into a set $\widetilde{S}=LS$ of rational points of height $O(N^{2})$ via an appropriate invertible linear transformation $L$ of $\mathbf{C}^{2}$ (see Lemma 2.1 below).
@@
    According to the above analysis, we would ideally like to show that if $S$ is not almost completely contained in a single line or circle, then $\widetilde{S}$ lies in the union of $O((\log N)^{O(1)})$ irreducible algebraic curves, defined over $\mathbf{Q}$, each of degree roughly $(\log N)^{C}$ (for some constant $C$).
@@
[…]
@@
    To tackle this issue, we begin by covering $\widetilde{S}^{\prime}$ by $O((\log N)^{O(1)})$ irreducible algebraic curves $\widetilde{\gamma}$, defined over $\mathbf{Q}$, of degree $O((\log N)^{C})$, contained in the surface $X_{k}$. The existence of such curves is ensured as a consequence of the determinant method from analytic number theory (see Theorem 2.2). Projecting the first two coordinates to $\mathbf{C}^{2}$, we are now closer to the aim in (2), but not quite there: we have covered $\widetilde{S}$ by an acceptable number of irreducible algebraic curves $\widetilde{\gamma}$ defined over $\mathbf{Q}$. However, these curves may have very low degree relative to the desired $(\log N)^{C}$.
@@
    We resolve this issue by utilizing the freedom we have to choose the points $P_{1},\ldots,P_{k}\in S$ we fix at the start of the argument. Each choice leads to a different $X_{k}$ with properties as in (4). We prove that, if $L^{-1}\widetilde{\gamma}$ is not a line or a circle, or if it is but there are enough points in $S$ outside it, then there is a choice of initial points $P_{1},\ldots,P_{k}$ such that the lift of $\widetilde{\gamma}$ on the corresponding $X_{k}$ is an irreducible curve of the correct degree $(\log N)^{C}$. Verifying that these lifts are irreducible is the crux of our argument.
@@ 2.1. Encoding the points of $S$ as rational points on a surface
We will use basic definitions and results from algebraic geometry throughout this paper; a good reference for this is [21, Chapter I]. Note that we may assume that $N$ is larger than any specified absolute constant, since the results are trivial for bounded $N$. In addition, observe that it suffices to prove Theorem 1.1 in the case that $S$ contains the origin and at least one other point. Indeed, either $S$ contains at least two points $P$ and $P^{\prime}$ in $N,N]^{2}$, or the result is trivial; moreover, if there exists a line or circle $C\subset\mathbf{R}^{2}$ such that all but $O((\log{2N})^{O(1)})$ points of $(S-P)\cap[-2N,2N]^{2}$ are contained in $C$, then certainly all but $O((\log{N})^{O(1)})$ points of $S$ are contained in a single line or circle. Thus, it follows from Lemma 2.1 that we may further assume, without loss of generality, that there exists a squarefree $m\in\mathbf{N}$ such that
@@
Let $k=k(N)$ be an integer parameter increasing slowly with $N$ (which we will, eventually, take to be $\asymp\log\log{N}$). If $|S|\leq k+1$, then we are done, since $N$ is large, by our initial assumption. Otherwise, pick $k$ distinct points $\tilde{P}_{j}=(a_{j},b_{j}\sqrt{m})\in S$ (so that $a_{i},b_{i}\in\frac{1}{2M}\mathbf{Z}\subseteq\mathbf{Q}$) for $1\leq j\leq k$, none of which is the origin, and set $P_{j}=(a_{j},b_{j})$ for each $1\leq j\leq k$. For any point $P=(a,b)\in\mathbf{Q}^{2}$, we define the polynomial $Q_{m,P}\in\mathbf{Q}[x,y,d]$ by
@@ 2.3. Counting rational points of small height on curves
$\mathcal{C}_{k}:=\left\{(x,y,d_{1},\dots,d_{k}):Q_{m,P_{j}^{\prime}}(x,y,d_{j})=0\text{ for all }j=1,\dots,k\right\}\cap(C^{\prime}\times\mathbf{C}^{k})\subset\mathbf{C}^{k+2}.$
Then, the points in $S\cap C$ correspond to a certain subset of rational points in the projective closure $\overline{\mathcal{C}_{k}}\subset\mathbf{P}^{k+2}$ of height at most $8N^{2}$. In Section 4, we will prove that $\overline{\mathcal{C}_{k}}$ is an irreducible curve of degree between $2^{k}$ and $2^{k}\deg{C}$, provided that $\tilde{P}_{1}^{\prime},\dots,\tilde{P}_{k}^{\prime}\in S$ are chosen appropriately. As was mentioned in the introduction, proving the irreducibility of $\overline{\mathcal{C}_{k}}$ is the most difficult part of the argument. In contrast to the proof of the irreducibility of $\overline{X_{k}}$, which is quite straightforward, the proof of the irreducibility of $\overline{\mathcal{C}_{k}}$ requires the normalization theorem for irreducible algebraic curves and the classification of compact Riemann surfaces combined with a detailed analysis of the geometry of the intersection of $C^{\prime}$ with a certain family of lines depending on the choice of the points $\tilde{P}_{1}^{\prime},\dots,\tilde{P}_{k}^{\prime}$. We emphasize that these points will be chosen on $C$ when $C$ is not a line or a circle, and outside of $C$ otherwise.
We can then apply a refinement of the Bombieri–Pila [5] bound due to Castryk, Cluckers, Dittmann, and Nguyen [7], who proved an upper bound on $|V(\mathbf{Q},N)|$, when $V\subset\mathbf{P}^{n}$ is an irreducible algebraic curve defined over $\mathbf{Q}$, with optimal dependence on $N$ and with an explicit polynomial dependence on the degree of $V$. By tracing through the proof of [7, Theorem 2] (and the results used to deduce it from the $n=2$ case [7, Proposition 4.3.2 and Lemma 5.1]), one can see that the dependence of their implied constant on $n$ is $\ll n^{200n^{2}/d}d^{4n^{2}/d}$. We record this in the following theorem.
@@
Thus, as long as $C$ either is not a line or circle or is a line or circle for which $|S\setminus C|\gg(\log\log{N})^{2}$, we can use Theorem 2.3 to bound the number of points in $S\cap C$ by $\ll e^{O(k)}N^{2/2^{k}}$, which, for $k\asymp\log\log N$, is $\ll(\log N)^{O(1)}$. Putting everything together yields Theorem 1.1. We will carry out the argument just outlined in Sections 34, and 6, and prove Propositions 1.5 and 1.6 in Section 5.
@@
Some polynomial dependence on $d$ is necessary in the bounds in Theorems  2.2 and 2.3 (see [7, Proposition 5]). As a consequence, no bound better than $|S\setminus C|\ll(\log{N})^{O(1)}$ in Theorem 1.1 seems attainable via our methods. The determinant method produces bounds that apply to any irreducible surface (respectively, curve) in $\mathbf{P}^{n}$ of degree $d$, but significantly better bounds could hold for individual surfaces (respectively, curves) in this collection. Thus, a further analysis of the specific varieties arising in our argument would be needed to improve Theorem 1.1.